please help?

1. Solve by graphing: x2 + y2 = 64 y = x + 8 . (x, y) = (smaller x-value) (x, y) = (larger x-value)
Answer #1

ny -y=x+8 1018) (-8,0) > x (8, 2 ty²=64 (0-8) Since from graph, we see that both x² + y ²64 4 nad y=x+d cuts at two points. T